OpenStudy (anonymous):

Chapter 5:Indices and Logarithms @ganeshie8

OpenStudy (anonymous):

\[\frac{ 49^{2n+1}\times14^{6-2n} }{ 56^{2n-5} }\]

OpenStudy (anonymous):

Simplify the following

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

start by factoring 14 and 56

OpenStudy (anonymous):

\[==\frac{ 49^{2n+1}\times(7\times2)^{6-2n} }{ (7\times8)^{2n-5} }\]

OpenStudy (anonymous):

Do it with 49 as well. You may want to factor all bases down to primes...

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[=\frac{ (7\times7)^{2n+1}\times(7\times2)^{6-2n} }{ (7\times8)^{2n-5} }\]

OpenStudy (anonymous):

\[=\frac{ (7\times7)^{2n+1}\times(7\times2)^{6-2n} }{ (7\times2\times2\times2)^{2n-5} }\]

OpenStudy (anonymous):

\[=\frac{ 7^{2n+1}\times7^{2n+1}\times7^{6-2n}\times2^{6-2n} }{ 7^{2n-5}\times2^{2n-5}\times2^{2n-5}\times2^{2n-5} }\]

OpenStudy (anonymous):

\[=\frac{ 7^{2n+1+2n+1+6-2n}\times2^{6-2n} }{ 7^{2n-5}\times2^{8n-15} }\]

OpenStudy (anonymous):

\[=\frac{ 7^{2n+8}\times2^{6-2n} }{ 7^{2n-5}\times2^{8n-15} }\]

OpenStudy (anonymous):

\[=7^{2n+8}\times2^{6-2n}\div7^{2n-5}\times2^{8n-15}\]

OpenStudy (anonymous):

\[=7^{2n+8-2n+5}\times2^{6-2n-8n+15}\]

OpenStudy (anonymous):

\[=7^{13}\times2^{21-10n}\]

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

I got \[=7^{13}\times2^{21-10n}\] but the answer in the book say it's \[=7^{13}\times2^{21-8n}\]

OpenStudy (anonymous):

Am i correct?

OpenStudy (unklerhaukus):

book is right

OpenStudy (unklerhaukus):

\[\large{ \frac{ 49^{2n+1}\times14^{6-2n} }{ 56^{2n-5} }\\\\ =\frac{7^{2(2n+1)}\times7^{6-2n}\times2^{6-2n}}{7^{2n-5}\times8^{2n-5}}\\\\ =\frac{7^{4n+2+6-2n}\times2^{6-2n}}{7^{2n-5}\times2^{3(2n-5)}}\\\\ =\frac{7^{2n+8}\times2^{6-2n}}{7^{2n-5}\times2^{6n-15)}}\\\\ =7^{2n+8-2n+5}\times2^{6-2n-6n+15}\\ =7^{13}\times 2^{21-8n} }\]

OpenStudy (unklerhaukus):

an \(\color{red}{\text e}\)rror in your working \[=\frac{ 7^{2n+1}\times7^{2n+1}\times7^{6-2n}\times2^{6-2n} }{ 7^{2n-5}\times2^{2n-5}\times2^{2n-5}\times2^{2n-5} }\]\[=\frac{ 7^{2n+1+2n+1+6-2n}\times2^{6-2n} }{ 7^{2n-5}\times2^{\color{red}6n-15} }\]

OpenStudy (anonymous):

okay know i get :) yay

OpenStudy (anonymous):

Thnx @UnkleRhaukus